A 12 kg box is resting on a table 1.5 m off the ground. It is then lifted up to 3.2 m off the ground. What is its increase in potential energy? (Round to 10 m/s²)




1
Expert's answer
2021-02-05T17:06:14-0500

"\Delta PE=mg\Delta h=mg(h_2-h_1)=12\cdot10\cdot(3.2-1.5)=204\ (J)" . Answer